Fermi National Laboratory


Accelerator Update

Wednesday July 7
The day shift began with much of the accelerator complex in access.

The Cryo Coordinator reported at 8:30 AM, that he thought the Tevatron (TeV) sector D1 quench might have been caused by a leaking kautzky valve that froze a vacuum o-ring, which caused a vacuum leak.

Linac technicians reported at 10:06 AM that they were replacing the cavity water pump for RF station #5 (LRF5).

Operations established beam to the Linac dump at 12:27 PM.

TeV turn on began at 12:46 PM.

Operations put the TeV through a dry squeeze at 2:49 PM.

Operations established beam in Booster at 3:05 PM.

Operations established beam to the Main Injector (MI) at 3:23 PM.

Operations established beam to MiniBooNE at 3:34 PM.

At 4:42 PM, Operations reported trying to turn on Switchyard (SY), but had trouble with a power supply (HP3US). They called in an expert.

An RF expert reported at 4:43 PM that all eighteen MI RF stations were operational.

Operations established beam to the Meson MTest experiment at 7:40 PM. The experiment requested fast spill.

Operations reported at 8:53 PM that after experts solved a number of problems the new Antiproton Source (Pbar) target became operational. Stacking resumed at 11:44 PM.

Thursday July 8
The midnight shift began with Pbar stacking, with Operations changing Switchyard (SY) beam from fast spill to slow spill, and with MiniBooNE and MTest taking beam.

Operations established beam to the Meson MCenter experiment around 12:30 AM.

At 2:12 AM, Operations reported that they were having trouble getting MI up to full intensity due to several RF stations (MIRF1, 8, 17) tripping off. Operators contacted an expert who had them make some adjustments.

Operations loaded the TeV with a 12x0 study-store at 3:11 AM.

At 4:51 AM, the Pbar lithium lens tripped off due to a water return pressure problem. It reset, but then tripped off again at 5:17 AM. An expert came in and bypassed this bit at 8:30 AM.

Operations put the TeV into shot setup at 5:12 AM.

At 6:16 AM, the TeV suffered a sector F2 & F3 quench.

The TeV shot setup resumed at 8:05 AM.

Operations established store 3629 at 9:43 AM with an initial luminosity of 56.4E30.

MCenter experimenters said that they were finished with beam at 11 AM.

The Cryo Coordinator reported at 11:26 AM that a TeV cryo compressor was off due to an oil leak. Cryo system techs will turn on another compressor after this store ends.

At 1:52 PM, the TeV suffered a sector F4 quench possibly due to a corrector magnet power supply (HF49) tripping off. A number of other houses also quenched.

Operators prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) for a Collision hall access at 2:33 PM. They were both out of access by 4:28 PM.

At 5 PM, a MI RF station's (MIRF1) gate bias switched off for the first of three times in thirty-four minutes. Operators reset the trips.

At 5:21 PM, A Linac RF station (KRF5) tripped off due to a gun spark. It reset okay.

TeV experts requested that operators replace the regulator and control cards for HF48 and HF49. A controls expert helped with this replacement.

At 5:43 PM, two MI RF stations (MIRF1 & 17) tripped off. They reset okay.

At 5:56 PM, two Booster power supplies (EAPS & WAPS) tripped off. The 400 MeV chopper also tripped. These trips all occurred during a timeline change. Everything reset okay.

At 6:33 PM, two MI RF stations (MIRF1 & 17) tripped off. They reset okay.

At 6:55 PM, KRF5 tripped off due to a gun spark for the first of two times in two minutes. It reset okay. It tripped off again about forty minutes later.

At 8:03 PM, Operations reported that the interlock system for Meson enclosure M01 dropped. Operators completed searching and securing the enclosure by 9:01 PM.

At 8:26 PM, the TeV suffered a sector A1 quench. The cause seemed to be due to HF49 power supply. Operators changed out its control card.

Operations reported at 8:54 PM that HF49 was still stuck on its flattop value. Operators continued to investigate.

Operations established a wet squeeze at 10:32 PM.

Operations loaded the TeV with a 12x0 study-store at 11:03 PM.

At 11:54 PM, Operations discovered a problem with a sector A quadrupole power supply (AQ0).

Friday July 9
The midnight shift began with TeV experts conducting studies, with operators investigating a problem with AQ0, and with MiniBooNE and MTest taking beam.

Operations called in a power supply expert to help with the AQ0 problems. He arrived at 2 AM and then called in another expert to help. They repaired the transducer electronics and had the supply online by 6:40 AM.

Operations put the TeV into shot setup at 7:51 AM.

Plans for Friday Morning and the Weekend
The plans for this morning include establishing a new store and then letting Pbar experts access their tunnels to conduct equipment maintenance and upgrades. Plans for this weekend include stacking, storing, and stashing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/9/2004   email Fermilab